use elimination method to solve.
-x+6y=-14
3x-6y=14

Answer by Mathtut(3670)

-x+6y=-14
3x-6y=14
:
add the two equations together and the y terms will be eliminated:
:
+6y-6y=0.....we are left with 3x-x=0
:
2x=0
:
x=0
:
now take x's found value and plug it back into either equation
:
-0+6y=-14
:
6y=-14
:
y=-14/6=-7/3
:
(0,-7/3)